What is the monthly payment on the following $400,000 mortgages: 30 yr, 3.95% 30 yr, 10.49%...

  1. What is the monthly payment on the following $400,000 mortgages:
  1. 30 yr, 3.95%
  2. 30 yr, 10.49%
  3. 15 yr, 3.95%
  4. 15 yr 10.49%
Answer #1

(A/P, I,n) = i((1 + i)^n)/((1 + i)^n-1)

a. I = 3.95% = 3.95% / 12 = 0.329167% per month

t = 30 * 12 = 360

Monthly payment = 400000 * (A/P, 0.329167%,360)

= 400000 * 0.004745

= 1898.15

b.

I = 10.49% = 10.49% / 12 = 0.874167% per month

t = 30 * 12 = 360

Monthly payment = 400000 * (A/P, 0.874167%,360)

= 400000 * 0.009140

= 3655.97

c.

.I = 3.95% = 3.95% / 12 = 0.329167% per month

t = 15 * 12 = 180

Monthly payment = 400000 * (A/P, 0.329167%,180)

= 400000 * 0.007372

= 2948.74

d.

I = 10.49% = 10.49% / 12 = 0.874167% per month

t = 15 * 12 = 180

Monthly payment = 400000 * (A/P, 0.874167%,180)

= 400000 * 0.011048

= 4419.12